Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 700,000 shares of common stock were outstanding. The interest rate on the bond payable was 10%, the income tax rate was 40%, and the dividend per share of common stock was $0.75 last year and $0.40 this year. The market value of the company’s common stock at the end of the year was $20. All of the company’s sales are on account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 1,100 $ 1,410
Accounts receivable, net 9,800 8,400
Inventory 13,600 11,200
Prepaid expenses 650 630
Total current assets 25,150 21,640
Property and equipment:
Land 10,700 10,700
Buildings and equipment, net 47,889 42,067
Total property and equipment 58,589 52,767
Total assets $ 83,739 $ 74,407
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 20,000 $ 19,200
Accrued liabilities 940 730
Notes payable, short term 0 290
Total current liabilities 20,940 20,220
Long-term liabilities:
Bonds payable 8,800 8,800
Total liabilities 29,740 29,020
Stockholders' equity:
Common stock 2,000 2,000
Additional paid-in capital 4,000 4,000
Total paid-in capital 6,000 6,000
Retained earnings 47,999 39,387
Total stockholders' equity 53,999 45,387
Total liabilities and stockholders' equity $ 83,739 $ 74,407
Weller Corporation
Comparative Income Statement and Reconciliation
(dollars in thousands)
This Year Last Year
Sales $ 68,000 $ 65,000
Cost of goods sold 34,000 32,000
Gross margin 34,000 33,000
Selling and administrative expenses:
Selling expenses 11,200 10,500
Administrative expenses 7,100 6,500
Total selling and administrative expenses 18,300 17,000
Net operating income 15,700 16,000
Interest expense 880 880
Net income before taxes 14,820 15,120
Income taxes 5,928 6,048
Net income 8,892 9,072
Dividends to common stockholders 280 525
Net income added to retained earnings 8,612 8,547
Beginning retained earnings 39,387 30,840
Ending retained earnings $ 47,999 $ 39,387

Required:

Compute the following financial ratios for this year:

1. Times interest earned ratio.

2. Debt-to-equity ratio.

3. Equity multiplier.

(For all requirements, round your answers to 2 decimal places.)
Answer #1

Following are the financial ratios for this year:

1. Times interest earned ratio:

= Income before Interest and taxes (EBIT) /  Interest Expenses

= 15700/880

= 17.84 times

2. Debt-to-equity ratio :

= Debt/ Equity

= 8800/ 6000+47999

=0.162

3. Equity multiplier :

= Total assets / Total stockholder's equity

= 83739/53999

= 1.55